# NUBO: A Transparent Python Package for Bayesian Optimization

### v2

@@ 2.1 Gaussian processes
A popular choice for the surrogate model $\mathcal{M}$ that acts as a representation of the objective function $f(\boldsymbol{x})$ is a Gaussian process [28, 30], a flexible non-parametric regression model. A Gaussian process is a finite collection of random variables that has a joint Gaussian distribution and is defined by a prior mean function $\mu_{0}(\boldsymbol{x}):\mathcal{X}\mapsto\mathbb{R}$ and a prior covariance kernel $\Sigma_{0}(\boldsymbol{x},\boldsymbol{x}^{\prime}):\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ resulting in the prior distribution
$f(\boldsymbol{X}_{n})\sim\mathcal{N}(m(\boldsymbol{X}_{n}),K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})),$ (3)
@@
Popular choices for the prior mean function $\mu_{0}(\cdot)$ are the zero and constant mean functions given in Equations 4 and 5. For the prior covariance function $\Sigma_{0}(\cdot,\cdot)$, the squared exponential kernel, also called the radial basis function (RBF) kernel, and the Matérn $\frac{5}{2}$ kernel are popular options. The covariance kernels in Equations 6 and 7 are based on the distance $r=\rvert\boldsymbol{x}-\boldsymbol{x}^{\prime}\lvert$ and have two parameters: the signal variance $\sigma^{2}_{f}$, sometimes also referred to as the output-scale, and the characteristic length-scale $l$. The former will scale the function with larger values, resulting in a more significant deviation from its mean, while the latter indicates how long function values are correlated along the input axes; the smaller the length-scale $l$, the shorter the correlation [28, 30].
@@
Covariance functions can be extended to include one characteristic length-scale $l_{d}$ for each input dimension $d$. In this case, input dimensions with large length scales are correlated for longer distances and are less relevant for changes in the prediction. This means that varying the values of the input dimension affects the prediction little. Input dimensions with small length scales are correlated for shorter distances, and even small changes in the input values can affect the prediction significantly. Gaussian processes with covariance functions that include multiple length scales are characterized by automatic relevance determination (ARD) of the input dimensions [31]. Here, the inverse of the length scales can be interpreted as the relevance of the corresponding dimensions [30]. The Gaussian process will estimate large length scales for irrelevant dimensions, automatically assigning them less importance.
@@
$f(\boldsymbol{X}_{*})\mid\mathcal{D}_{n},\boldsymbol{X}_{*}\sim\mathcal{N}\left(\mu_{n}(\boldsymbol{X}_{*}),\sigma^{2}_{n}(\boldsymbol{X}_{*})\right)$ (8)
$\mu_{n}(\boldsymbol{X}_{*})=K(\boldsymbol{X}_{*},\boldsymbol{X}_{n})\left[K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})+\sigma^{2}_{y}I\right]^{-1}(\boldsymbol{y}-m(\boldsymbol{X}_{n}))+m(\boldsymbol{X}_{*})$ (9)
$\sigma^{2}_{n}(\boldsymbol{X}_{*})=K(\boldsymbol{X}_{*},\boldsymbol{X}_{*})-K(\boldsymbol{X}_{*},\boldsymbol{X}_{n})\left[K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})+\sigma^{2}_{y}I\right]^{-1}K(\boldsymbol{X}_{n},\boldsymbol{X}_{*}),$ (10)
@@
The hyper-parameters $\theta$ of the Gaussian process, for example, the constant $c$ in the mean function, the signal variance $\sigma^{2}_{f}$ and characteristic length-scales $\boldsymbol{l}$ in the covariance kernel, and the noise variance $\sigma^{2}_{y}$, can be estimated by maximizing the log-marginal likelihood in Equation 11 via maximum likelihood estimation (MLE) [30].
@@ 2.2 Acquisition functions
Acquisition functions use the posterior distribution of the Gaussian process to compute a criterion that assesses if a test point is a good potential candidate point to evaluate via the objective function $f(\boldsymbol{x})$. Thus, maximizing the acquisition function suggests the test point that, based on the current training data $\mathcal{D}_{n}$, has the highest potential and information gain to get closer to the global optimum while exploring the input space. To do this, an acquisition function $\alpha(\cdot)$ balances exploration and exploitation. The former is characterized by areas with no or only a few observed data points where the uncertainty of the Gaussian process is high and the latter by areas where the posterior mean of the Gaussian process is large. This exploration-exploitation trade-off ensures that Bayesian optimization does not converge to the first (potentially local) maximum it encounters but efficiently explores the entire input space.
@@ 2.2.1 Analytical acquisition functions
NUBO supports two of the most popular acquisition functions whose performance has been demonstrated in both theoretical and empirical research. Expected improvement (EI) [9] selects points with the biggest potential to improve on the current best observation, while upper confidence bound (UCB) [32] takes an optimistic view of the posterior uncertainty and assumes it to be true to a user-defined level. Expected improvement (EI) is defined as
@@
where $\beta$ is a predefined trade-off parameter, and $\mu_{n}(\cdot)$ and $\sigma_{n}(\cdot)$ are the mean and the standard deviation of the posterior distribution of the Gaussian process. For guidance on the choice of $\beta$, consult the theoretical properties in [32] or empirical conclusions in [12].
Both of these acquisition functions can be maximized with a deterministic optimizer, such as L-BFGS-B [3] for bounded unconstrained problems or SLSQP [33] for bounded constrained problems. However, the use of analytical acquisition functions is restricted to sequential single-point problems for which every point suggested by Bayesian optimization is observed via the objective function $f(\boldsymbol{x})$ immediately before the optimization loop is repeated.
@@ 2.2.2 Monte Carlo acquisition functions
For parallel multi-point batches or asynchronous optimization, the analytical acquisition functions are generally intractable. To use Bayesian optimization in these cases, NUBO supports approximating the analytical acquisition function through Monte Carlo sampling [10, 27].
The idea is to draw many samples directly from the posterior distribution and then approximate the acquisition functions by averaging these Monte Carlo samples. This method is made viable by reparameterizing the acquisition functions and the computation of samples from the posterior distribution via base samples randomly drawn from a standard normal distribution $z\sim\mathcal{N}(0,1)$. Thus, the analytical acquisition functions from in Equations 12 and 13 can be approximated as
$\alpha_{\text{EI}}^{\text{MC}}(\boldsymbol{X}_{*})=\max\left(ReLU(\mu_{n}(\boldsymbol{X}_{*})+\boldsymbol{L}\boldsymbol{z}-y^{best})\right)$ (14)
$\alpha_{\text{UCB}}^{\text{MC}}(\boldsymbol{X}_{*})=\max\left(\mu_{n}(\boldsymbol{X}_{*})+\sqrt{\frac{\beta\pi}{2}}\lvert\boldsymbol{L}\boldsymbol{z}\rvert\right),$ (15)
@@
Due to the randomness in the Monte Carlo samples, these acquisition functions can only be optimized by stochastic optimizers, such as Adam [2]. However, there is some empirical evidence that fixing the base samples for individual Bayesian optimization loops does not affect the performance negatively [22]. This method would allow deterministic optimizers, such as L-BFGS-B [3] and SLSQP [33], to be used but could potentially introduce bias due to sampling randomness.
Two optimization strategies for multi-point batches are proposed in the literature [27]: The first is a joint optimization approach, where the acquisition functions are optimized over all batch points simultaneously. The second option is a greedy sequential approach where one point after another is selected, holding all previous points fixed until the batch is complete. Empirical evidence shows that both methods approximate the acquisition successfully. However, the greedy approach seems to have a slight edge over the joint strategy for some examples [27]. It is also faster to compute for larger batches.
Asynchronous optimization [10] leverages the same property as sequential greedy optimization: the pending points that have not yet been evaluated can be added to the test points but are treated as fixed. In this way, they affect the joint multivariate normal distribution but are not considered directly in the optimization. Asynchronous optimization is particularly beneficial for objective functions for which the evaluation time varies. In these cases, the optimization can be continued while some points are still being evaluated.
@@ 3 NUBO
NUBO is a Bayesian optimization package in Python that focuses on transparency and user experience to make Bayesian optimization accessible to researchers from a wide range of disciplines whose expertise is not necessarily statistics or computer science. With this overall goal in mind, NUBO ensures transparency by implementing clean and comprehensible code, precise references, and thorough documentation within this research article and on our website at www.nubopy.com. We avoid implementing overly complex and convoluted functions and objects that require the retracing of individual elements through multiple files to be fully understood. We prioritize user experience defined by a modular and flexible design that can be intuitively tailored to unique problems, easy-to-read and write syntax, and a careful selection of Bayesian optimization algorithms. The latter is essential as we try not to overwhelm the user with a larger number of options but instead focus on what is essential to successfully optimize computer simulators and physical experiments.
To create a powerful package with good longevity, starting with a strong foundation is important. NUBO is built upon the Torch [^4] See https://pytorch.org/ for documentation and https://pypi.org/project/torch/ for package information on PyPI. ecosystem [34] that provides a strong scientific computation framework for working with tensors, a selection of powerful optimization algorithms, such as torch.Adam [2], automatic differentiation capabilities to compute gradients of acquisition functions via torch.autograd, and GPU acceleration. Furthermore, GPyTorch [^5] See https://gpytorch.ai/ for documentation and https://pypi.org/project/gpytorch/ for package information on PyPI. [35], the package we use to implement Gaussian processes for our surrogate modeling, is also based in Torch and combines seamlessly with NUBO. We borrow the L-BFGS-B [3] and SLSQP [33] optimization algorithms from SciPy [^6] See https://scipy.org/ for documentation and https://pypi.org/project/scipy/ for package information on PyPI. [36] for the deterministic optimization of the acquisition functions and use NumPy [^7] See https://numpy.org/ for documentation and https://pypi.org/project/numpy/ for package information on PyPI. [37] to make data suitable for these optimizers.
NUBO and all its required dependencies can be installed from the Python Package Index (PyPI) [38] with the packet installer pip [39] via the terminal:
@@ 3.1 Gaussian processes
NUBO uses the GPyTorch [35] package to implement Gaussian processes for surrogate modeling. While GPyTorch allows the definition of many different Gaussian processes through its various mean functions, covariance kernels, and methods for hyper-parameter estimation, we provide a predefined Gaussian process in the nubo.models module that follows the work of [10]. The GaussianProcess is specified by a constant mean function, and the Matérn $\frac{5}{2}$ ARD kernel that, due to its flexibility, is well suited for practical optimization as it can represent a wide variety of real-world objective functions. The code below implements a Gaussian process and estimates its hyper-parameters from some training inputs x_train and training outputs y_train by maximizing the log-marginal likelihood in Equation 11 with the fit_gp function. The hyper-parameters include the constant in the mean function, the output scale and length scales in the covariance kernel, and the noise in the Gaussian likelihood. The training inputs and training outputs are specified as a torch.Tensor of size $n\times d$ and length $n$, respectively, where $n$ is the number of points, and $d$ is the number of input dimensions. Calling the function fit_gp results in a trained Gaussian process that can subsequently be used for Bayesian optimization.
@@
    $>$ from nubo.models import GaussianProcess, fit_gp    $>$ from gpytorch.likelihoods import GaussianLikelihood    $>$    $>$    $>$ likelihood = GaussianLikelihood()    $>$ gp = GaussianProcess(x_train, y_train, likelihood=likelihood)    $>$ fit_gp(x_train, y_train, gp=gp, likelihood=likelihood)While Gaussian processes can estimate noise, for example, observational noise occurring when taking measurements from the data, we might prefer specifying the noise explicitly if it is known. We can exchange the GaussianLikelihood for the FixedNoiseGaussianLikelihood and specify the noise for each training point. The FixedNoiseGaussianLikelihood allows us to decide if any additional noise should be estimated by setting the learn_additional_noise attribute to True or False. The snippet below fixes the observational noise of each training point at 2.5% and estimates any additional noise.
    $>$ from nubo.models import GaussianProcess, fit_gp    $>$ from gpytorch.likelihoods import FixedNoiseGaussianLikelihood    $>$    $>$    $>$ noise = torch.ones(x_train.size(0)) * 0.025    $>$ likelihood = FixedNoiseGaussianLikelihood(noise=noise,    ...                                           learn_additional_noise=True)    $>$ gp = GaussianProcess(x_train, y_train, likelihood=likelihood)    $>$ fit_gp(x_train, y_train, gp=gp, likelihood=likelihood)
3.2 Bayesian optimization

### v3

@@ 2.1 Gaussian processes
A popular choice for the surrogate model $\mathcal{M}$ that acts as a representation of the objective function $f(\boldsymbol{x})$ is a Gaussian process (Gramacy, 2020; Rasmussen and Williams, 2006), a flexible non-parametric regression model. A Gaussian process is a finite collection of random variables that has a joint Gaussian distribution and is defined by a prior mean function $\mu_{0}(\boldsymbol{x}):\mathcal{X}\mapsto\mathbb{R}$ and a prior covariance kernel $\Sigma_{0}(\boldsymbol{x},\boldsymbol{x}^{\prime}):\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ resulting in the prior distribution
$f(\boldsymbol{X}_{n})\sim\mathcal{N}(m(\boldsymbol{X}_{n}),K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})),$
@@
Popular choices for the prior mean function $\mu_{0}(\cdot)$ are the zero and constant mean functions given in Equations 2 and 3. For the prior covariance function $\Sigma_{0}(\cdot,\cdot)$, the squared exponential kernel, also called the radial basis function (RBF) kernel, and the Matérn $\frac{5}{2}$ kernel are popular options. The covariance kernels in Equations 4 and 5 are based on the distance $r=\rvert\boldsymbol{x}-\boldsymbol{x}^{\prime}\lvert$ and have two parameters: the signal variance $\sigma^{2}_{f}$, sometimes also referred to as the output-scale, and the characteristic length-scale $l$. The former will scale the function with larger values resulting in a larger deviation from its mean, while the latter indicates for how long function values are correlated along the input axes, the smaller the length-scale $l$ the shorter the correlation (Gramacy, 2020; Rasmussen and Williams, 2006).
@@
Covariance functions can be extended to include one characteristic length-scale $l_{d}$ for each input dimension $d$. In this case, input dimensions with large length-scales are correlated for longer distances and are less relevant for changes in the prediction. This means that varying the values of the input dimension affects the prediction little. Input dimensions with small length-scales are correlated for shorter distances, and even small changes in the input values can affect the prediction significantly. Gaussian processes with covariance functions that include multiple length-scales are characterised by automatic relevance determination (ARD) of the input dimensions (Neal, 1996). Here, the inverse of the length-scales can be interpreted as the relevance of the corresponding dimensions (Rasmussen and Williams, 2006). The Gaussian process will estimate large length-scales for irrelevant dimensions, automatically assigning them less importance.
@@
$f(\boldsymbol{X}_{*})\mid\mathcal{D}_{n},\boldsymbol{X}_{*}\sim\mathcal{N}\left(\mu_{n}(\boldsymbol{X}_{*}),\sigma^{2}_{n}(\boldsymbol{X}_{*})\right)$
$\mu_{n}(\boldsymbol{X}_{*})=K(\boldsymbol{X}_{*},\boldsymbol{X}_{n})\left[K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})+\sigma^{2}_{y}I\right]^{-1}(\boldsymbol{y}-m(\boldsymbol{X}_{n}))+m(\boldsymbol{X}_{*})$
$\sigma^{2}_{n}(\boldsymbol{X}_{*})=K(\boldsymbol{X}_{*},\boldsymbol{X}_{*})-K(\boldsymbol{X}_{*},\boldsymbol{X}_{n})\left[K(\boldsymbol{X}_{n},\boldsymbol{X}_{n})+\sigma^{2}_{y}I\right]^{-1}K(\boldsymbol{X}_{n},\boldsymbol{X}_{*}),$
@@
The hyper-parameters $\theta$ of the Gaussian process, for example, the constant $c$ in the mean function, the signal variance $\sigma^{2}_{f}$ and characteristic length-scales $\boldsymbol{l}$ in the covariance kernel, and the noise variance $\sigma^{2}_{y}$, can be estimated by maximising the log-marginal likelihood in Equation 6 via maximum likelihood estimation (MLE) (Rasmussen and Williams, 2006).
@@ 2.2 Acquisition functions
Acquisition functions use the posterior distribution of the Gaussian process to compute a criterion that assesses if a test point is a good potential candidate point to evaluate via the objective function $f(\boldsymbol{x})$. Thus, maximising the acquisition function suggests the test point that, based on the current training data $\mathcal{D}_{n}$, has the highest potential and information gain to get closer to the global optimum while exploring the input space. To do this, an acquisition function $\alpha(\cdot)$ balances exploration and exploitation. The former is characterised by areas with no or only a few observed data points where the uncertainty of the Gaussian process is high, and the latter by areas where the posterior mean of the Gaussian process is large. This exploration-exploitation trade-off ensures that Bayesian optimisation does not converge to the first (potentially local) maximum it encounters but efficiently explores the full input space.
@@ 2.2.1 Analytical acquisition functions
NUBO supports two of the most popular acquisition functions whose performance has been demonstrated in both theoretical and empirical research. Expected improvement (EI) (Jones et al., 1998) selects points with the biggest potential to improve on the current best observation, while upper confidence bound (UCB) (Srinivas et al., 2010) takes an optimistic view of the posterior uncertainty and assumes it to be true to a user-defined level. Expected improvement (EI) is defined as
@@
where $\beta$ is a predefined trade-off parameter, and $\mu_{n}(\cdot)$ and $\sigma_{n}(\cdot)$ are the mean and the standard deviation of the posterior distribution of the Gaussian process. For guidance on the choice of $\beta$, consult the theoretical properties in Srinivas et al. (2010) or empirical conclusions in Diessner et al. (2022).
Both of these acquisition functions can be maximised with a deterministic optimiser, such as L-BFGS-B (Zhu et al., 1997) for bounded unconstrained problems or SLSQP (Kraft, 1994) for bounded constrained problems. However, the use of analytical acquisition functions is restricted to sequential single-point problems for which every point suggested by Bayesian optimisation is observed via the objective function $f(\boldsymbol{x})$ immediately before the optimisation loop is repeated.
@@ 2.2.2 Monte Carlo acquisition functions
For parallel multi-point batches or asynchronous optimisation, the analytical acquisition functions are, in general, intractable. To use Bayesian optimisation in these cases, NUBO supports the approximation of the analytical acquisition function through Monte Carlo sampling (Snoek et al., 2012; Wilson et al., 2018).
The idea is to draw a large number of samples directly from the posterior distribution and then approximate the acquisition functions by averaging these Monte Carlo samples. This method is made viable by the reparameterisation of the acquisition functions and the computation of samples from the posterior distribution via base samples randomly drawn from a standard normal distribution $z\sim\mathcal{N}(0,1)$. Thus, the analytical acquisition functions from in Equations 7 and 8 can be approximated as
$\alpha_{\text{EI}}^{\text{MC}}(\boldsymbol{X}_{*})=\max\left(ReLU(\mu_{n}(\boldsymbol{X}_{*})+\boldsymbol{L}\boldsymbol{z}-y^{best})\right)$
[…]
@@
Due to the randomness in the Monte Carlo samples, these acquisition functions can only be optimised by stochastic optimisers, such as Adam (Kingma and Ba, 2014). However, there is some empirical evidence that fixing the base samples for individual Bayesian optimisation loops does not affect the performance negatively (Balandat et al., 2020). This method would allow deterministic optimisers, such as L-BFGS-B (Zhu et al., 1997) and SLSQP (Kraft, 1994), to be used but could potentially introduce bias due to sampling randomness.
Two optimisation strategies for multi-point batches are proposed in the literature (Wilson et al., 2018): The first is a joint optimisation approach, where the acquisition functions are optimised over all points of the batch simultaneously. The second option is a greedy sequential approach where one point after another is selected, holding all previous points fixed until the batch is full. Empirical evidence shows that both methods approximate the acquisition successfully. However, the greedy approach seems to have a slight edge over the joint strategy for some examples (Wilson et al., 2018). It is also faster to compute for larger batches.
Asynchronous optimisation (Snoek et al., 2012) leverages the same property as sequential greedy optimisation: the pending points that have not yet been evaluated can be added to the test points but are treated as fixed. In this way, they affect the joint multivariate normal distribution but are not considered directly in the optimisation. Asynchronous optimisation is particularly beneficial for objective functions for which the evaluation time varies. In these cases, the optimisation can be continued while some points are still being evaluated.
@@ 3 NUBO
NUBO is a Bayesian optimisation package in Python that focuses on transparency and user experience to make Bayesian optimisation accessible to researchers from a wide range of disciplines whose area of expertise is not necessarily statistics or computer science. With this overall goal in mind, NUBO ensures transparency by implementing clean and comprehensible code, precise references and thorough documentation within this research article and on our website at www.nubopy.com. We avoid implementations of overly complex and convoluted functions and objects that require the retracing of individual elements through multiple files to be fully understood. We prioritise user experience defined by a modular and flexible design that can be intuitively tailored to unique problems, easy-to-read and write syntax, and a careful selection of Bayesian optimisation algorithms. The latter is important as we try not to overwhelm the user with a larger number of options but rather focus on what is essential to optimise computer simulators and physical experiments successfully.
To create a powerful package with good longevity, it is important to start with a strong foundation. NUBO is built upon the Torch [^4] See https://pytorch.org/ for documentation and https://pypi.org/project/torch/ for package information on PyPI. ecosystem (Paszke et al., 2019) that provides a strong scientific computation framework for working with tensors, a selection of powerful optimisation algorithms, such as torch.Adam (Kingma and Ba, 2014), automatic differentiation capabilities to compute gradients of acquisition functions via torch.autograd, and GPU acceleration. Furthermore, GPyTorch [^5] See https://gpytorch.ai/ for documentation and https://pypi.org/project/gpytorch/ for package information on PyPI. (Gardner et al., 2018), the package we use to implement Gaussian processes for our surrogate modelling, is also based in Torch and combines seamlessly with NUBO. We borrow the L-BFGS-B (Zhu et al., 1997) and SLSQP (Kraft, 1994) optimisation algorithms from SciPy [^6] See https://scipy.org/ for documentation and https://pypi.org/project/scipy/ for package information on PyPI. (Virtanen et al., 2020) for the deterministic optimisation of the acquisition functions and use NumPy [^7] See https://numpy.org/ for documentation and https://pypi.org/project/numpy/ for package information on PyPI. (Harris et al., 2020) to make data suitable for these optimisers.
NUBO and all its required dependencies can be installed from the Python Package Index (PyPI) (Python Software Foundation, n.d.) with the packet installer pip (The pip developers, 2023) via the terminal:
@@ 3.1 Gaussian processes
NUBO uses the GPyTorch (Gardner et al., 2018) package to implement Gaussian processes for surrogate modelling. While GPyTorch allows the definition of many different Gaussian processes through its various mean functions, covariance kernels, and methods for hyper-parameter estimation, we provide a predefined Gaussian process in the nubo.models module that follows the work of Snoek et al. (2012). The GaussianProcess is specified by a constant mean function and the Matérn $\frac{5}{2}$ ARD kernel that, due to its flexibility, is well suited for practical optimisation as it can represent a wide variety of real-world objective functions. The code below implements a Gaussian process and estimates its hyper-parameters from some training inputs x_train and training outputs y_train by maximising the log-marginal likelihood in Equation 6 with the fit_gp function. The hyper-parameters include the constant in the mean function, the output-scale and length-scales in the covariance kernel, and the noise in the Gaussian likelihood. The training inputs and training outputs are specified as a torch.Tensor of size $n\times d$ and length $n$, respectively, where $n$ is the number of points and $d$ is the number of input dimensions. Calling the function fit_gp results in a trained Gaussian process that can subsequently be used for Bayesian optimisation.
    >>> from nubo.models import GaussianProcess, fit_gp
    >>> from gpytorch.likelihoods import GaussianLikelihood
    >>> likelihood = GaussianLikelihood()
    >>> gp = GaussianProcess(x_train, y_train, likelihood = likelihood)
    >>> fit_gp(x_train, y_train, gp = gp, likelihood = likelihood)
@@
While Gaussian processes are capable of estimating noise, for example, observational noise occurring when taking measurements from the data, we might prefer specifying the noise explicitly if it is known. In these cases, we can exchange the GaussianLikelihood for the FixedNoiseGaussianLikelihood and specify the noise for each training point. The FixedNoiseGaussianLikelihood allows us to decide if any additional noise should be estimated by setting the learn_additional_noise attribute to True or False. The snippet below fixes the observational noise of each training point at 2.5% and estimates any additional noise.
    >>> from nubo.models import GaussianProcess, fit_gp
    >>> from gpytorch.likelihoods import FixedNoiseGaussianLikelihood
    >>> noise = torch.ones(x_train.size(0)) * 0.025
    >>> likelihood = FixedNoiseGaussianLikelihood(noise = noise,
    ...                                           learn_additional_noise = True)
    >>> gp = GaussianProcess(x_train, y_train, likelihood = likelihood)
    >>> fit_gp(x_train, y_train, gp = gp, likelihood = likelihood)
3.2 Bayesian optimisation